OpenStudy (anonymous):

A jet plane is traveling with a velocity of 720 kilometers/hour due north. There is a crosswind of 40 kilometers/hour from the east. Find the resultant velocity and the direction of the jet plane.

OpenStudy (festinger):

 \[Tan\Theta=\frac{opposite}{adjacent}\] From the diagram, the triangle formed from the velocities makes the "opposite" 40 and "adjacent" 720. Thus, \[\Theta=3.17^{o} from\:the\:North\] Not that when reporting angles in bearing, the North is 0 and it sweeps clockwise. The magnitude of the resultant belocity is simply: \[\sqrt{720^{2}+40^{2}}=721km/h\] You can check if the angles are correct if you apply the relation: \[\frac{40}{720}=sin\Theta\] If the values are the same, it shows there there aren't any mistakes!